find the position of a vertex for the graph of the function f(x) = x to the second power+6x+7
am desperate!!

Answer by scott8148(6628)

the vertex is on the axis of symmetry

for a quadratic of the form ax^2+bx+c, the equation for the axis is x=-b/2a

x=-6/2 or x=-3

to find the y-coordinate, substitute into the equation __ y=(-3)^2+6(-3)+7 __ y=-2

so the vertex is (-3,-2)